# Sub-nanometer 3D morphometric precision of polarisation-resolved wide-field optical extinction microscopy determines the roundness of individual gold nanospheres

### v1

@@ Abstract
Quantitative polarisation-resolved optical extinction microscopy of individual plasmonic nanoparticles has recently been introduced as a powerful tool to characterise the nanoparticle’s morphology with a precision comparable to electron microscopy, while using a simple optical microscope [Nanoscale 12, 16215 (2020)]. Here we extend the technique by adding measurements for radial polarisation in the condenser back focal plane, probing plasmonic resonances polarised in axial direction. The combined linear and radial polarisation measurements provide a significantly enhanced precision of the retrieved 3D morphology, as we show on defect-free ultra-uniform gold nanospheres of 30 nm nominal diameter characterised by transmission electron microscopy. The measured cross-sections are quantitatively described by an ellipsoid model, determining the three semi-axes and rotation angles by fitting the measurements. Evaluation the distribution of the fit error across the set of measured particles, the material permittivity dataset and surface damping parameter $g$ providing the best fit are found to be the single crystal dataset by Olmon et al. [Phys. Rev. B 86, 235147 (2012)] and $g\approx 1.8$, respectively. The precision of the retrieved aspect ratio is below 5%, and all three ellipsoidal semi-axes are determined with an impressive precision of $\pm 0.25\,\mathrm{nm}$. Notably, corrections to the Rayleigh-Gans ellipsoid model due to retardation are significant even though the particle diameters are more than an order of magnitude smaller than the wavelength, and taking them into account improves the accuracy to below a nanometer.
@@ I Introduction
Recently, our group established a methodology for 3D morphometry of single NPs called the Optical Nanosizer ($\mathrm{ON}$).[43] As a wide-field microscopy modality, the ON offers simultaneous measurement of up to hundreds of NPs in a single field of view, with a sensitivity down to $\sigma_{\mathrm{ext}}\approx 1\,$nm2, similar to the cross-section of a gold nanosphere (GNS) of $D\approx 2\,$nm diameter. In our method, a series of brightfield transmission images of NPs deposited on a coverslip are captured on a conventional microscope under illumination at different wavelengths and linear polarisations. Dark-field images can be additionally taken to distinguish scattering and absorption cross-sections.[35] Images are analysed as differential transmission contrast, to quantify $\sigma_{\mathrm{ext}}$ for each NP in the field of view. By comparing the Rayleigh–Gans model for ellipsoidal particles with the measurements of $\sigma_{\mathrm{ext}}$ for each NP in the images, we inferred the size and shape of nominally spherical small (10 to 30 nm diameter) gold nanoparticles, and gold nanorods (10 nm diameter, 30 nm length). The precision of the method was found to be $\pm 0.5\,\mathrm{nm}$ for size, and about $10\%$ for the shape anisotropy. In terms of accuracy, we found the measurements of GNS having nominal diameter, D=30D=30\,nm, to be within 11\,nm of TEM characterization on the same sample, i.e. showing only about $4\%$ discrepancy. Notably, for shape retrieval we used linearly polarised excitation light in the condenser back focal plane with varying in-plane angle. Even with a high numerical aperture (NA) condenser focussing the light, only a small fraction of light at the sample is polarised out of plane, i.e. axially. Hence, the measurements are dominated by excitation polarisations in the sample plane, and thus are more sensitive to the in-plane geometry of the NP. The out-of-plane geometry was therefore indirectly retrieved by spectral properties using the dependence of surface plasmon resonance frequencies on particle morphology.
@@ II Model: Permittivity, polarisability, and optical cross-sections
The previous $\mathrm{ON}$ analytical developments were described in detail in Ref. 43, 44. We build here on this analysis, extending it to add the treatment of the radially polarised illumination. We briefly recap the model in the following. The morphometry of each particle is derived from the measured cross-sections by solving ‘the inverse problem,’ using a theoretical framework relating the geometry of a NP to its optical cross-sections. Modelling the optical cross-sections requires knowledge of the complex permittivity of the NP material. In the literature, different measured spectra of the permittivity of gold are provided. We considered three of these, namely the data of Johnson & Christy (JC),[45] McPeak et al. (MP),[46] and Olmon et al. (OL).[47] The measurements were made by spectroscopic ellipsometry on thin polycrystalline films (JC, MP) or single crystalline gold (OL). To account for a surface damping effect in the NPs, we use the following approach. The permittivity is described by a Drude model with a bound electron contribution stemming from the interband transitions from the $d$ bands into the conduction band, and surface damping effects, given by,[48, 35]
@@
We use the Rayleigh–Gans model, describing ellipsoidal NPs much smaller than the light wavelength ($\lambda$), to determine the NP polarisability using the complex material permittivity and geometry of the NP, and the permittivity, $\epsilon_{\mathrm{m}}$, of the surrounding medium. For consistency with Ref. 44, we maintain a Cartesian reference system of unit vectors $\vec{e}_{\kappa}^{\prime}$, with $\kappa=x^{\prime},\,y^{\prime}$, and $z^{\prime}$ which respectively point in the directions of the three semi-axes of an ellipsoidal NP, ($a$, $b$, $c$), which we order such that $a\geq b\geq c$. In this basis, the polarisability tensor $\alpha^{\prime}$ is a diagonal matrix. The individual entries are the polarisabilities for fields oriented along each of the NP’s semi-axes, with each given by
@@
where $V$ is the NP volume, the free space permittivity is $\epsilon_{0}$, the dielectric function of the surrounding medium is $\epsilon_{\mathrm{m}}$, and the depolarisation factors are $L_{j}$, with $j\in\{1,\,2,\,3\}$. The geometry of the NP determines $L_{j}$. For example, for a sphere, $L_{j}=1/3$. $\epsilon_{\mathrm{m}}$ is assumed to be real and frequency independent. All materials are assumed to be non-magnetic, i.e. with relative permeability $\mu_{\mathrm{r}}=1$. We note that metals exhibit a zero crossing versus wavelength of the real part of the denominator in Eq. 3 at $\Re\{\epsilon_{\mathrm{NP}}\}=(1-L_{j}^{-1})\epsilon_{\mathrm{m}}$, which results in a peak in $\sigma_{\mathrm{ext}}$ called the localised surface plasmon resonance (LSPR). For an incident field $\vec{E}^{\prime}$, the induced electric dipole moment is given by
@@
We choose a Cartesian coordinate system to describe the laboratory frame, with axes $\iota\in\{x,y,z\}$, and associated unit vectors $\vec{e}_{\iota}$, where $\vec{e}_{z}$ points along the optical path, and $\vec{e}_{x}$, $\vec{e}_{y}$ span the sample plane. To transform $\alpha^{\prime}$ into the laboratory reference frame. We define[43] the 3D rotation matrix using $R=R_{\psi}R_{\theta}R_{\phi}$ and $R^{\intercal}=R_{\phi}^{\intercal}R_{\theta}^{\intercal}R_{\psi}^{\intercal}$, with $\phi$, $\theta$, $\psi$ the angles of rotation about $\vec{e}_{x}$, $\vec{e}_{y}$, $\vec{e}_{z}$, respectively. We transform a vector from the NP frame into the laboratory frame with $\vec{v}=R\vec{v}^{\prime}$. Eq. 4 becomes $R^{\intercal}\vec{p}=\alpha^{\prime}R^{\intercal}\vec{E}$, such that the polarisability in the laboratory frame is given by
@@ III.3 Optical Extinction Measurements
In Ref. 53 we detailed our procedure to quantitatively measure $\sigma_{\mathrm{ext}}$, with additional information given in Ref. 54. In short, two brightfield images called $I_{1}$ and $I_{2}$ are captured, which differ only by a lateral shift of the sample, typically 1 or 2 µm (i.e. a few optical resolutions). $I_{1}$ and $I_{2}$ are averaged in real time over a number, $N_{\mathrm{i}}$, of individual acquisitions to reduce shot noise and data storage requirements. We compute $\Delta_{1,2}=1-I_{1,2}/I_{2,1}$ to yield two extinction images (with inverted contrast). To further reduce shot noise as well as systematic noise due to sensor electronic drift, we can average the extinction over a number of repetitions $N_{\mathrm{r}}$. For a NP centered in the areas $A_{1,2}$ of radius $r_{\mathrm{i}}$ in $I_{1,2}$, we determine its extinction as $2\sigma_{\mathrm{ext}}=\int_{A_{1}}\Delta_{1}\mathrm{d}A+\int_{A_{2}}\Delta_{2}\mathrm{d}A$. We calculate the shot-noise-limited standard deviation in the measured $\sigma_{\mathrm{ext}}$ as [53]
@@
where $M$ is the magnification from the sample to the detector, $N_{\mathrm{a}}=N_{\mathrm{i}}N_{\mathrm{r}}$ is the number of acquired frames, and $d_{\mathrm{px}}=6.5\,$µm is the pixel pitch of the sensor. We previously used[34] a measurement radius of $r_{\mathrm{i}}=3\lambda/(2\mathrm{NA})$ to analyse $\sigma_{\mathrm{ext}}$, and the effect of $r_{\mathrm{i}}$ on $\sigma_{\mathrm{ext}}$ is characterised in detail in Ref. 55. In the present work, we employ three values of $r_{\mathrm{i}}$, specifically $r_{\mathrm{s}}=3\lambda/(2\mathrm{NA})$, $r_{\mathrm{h}}=1.2\,r_{\mathrm{s}}$, and $r_{\mathrm{l}}=0.8\,r_{\mathrm{s}}$, with $N_{\mathrm{i}}=128$ and $N_{\mathrm{r}}=20$. At $\Lambda=550\,$nm, Eq. 8 yields $\hat{\sigma}_{\mathrm{ext}}\approx 7.5\,\mathrm{nm}^{2}$ using $r_{\mathrm{s}}$, while the noise seen in the extinction data at this wavelength is $15.6\,\mathrm{nm}^{2}$. The additional noise is attributed to surface roughness, debris, and/or residual sensor fluctuations, as analysed in detail in Ref. 53. We note that measurements are taken over several different fields of view (FOV) and aggregated for the final statistics. The $\hat{\sigma}_{\mathrm{ext}}$ reported are the polarisation-averaged values of the noise, averaged over the measured values from all the FOVs. We found that $\hat{\sigma}_{\mathrm{ext}}$ varies by $\pm 5\%$ across the illumination polarisations, and by $\pm 15\%$ across FOVs. We measure $\sigma_{\mathrm{ext}}$ for multiple wavelengths $\Lambda$, both for the linear polariser at multiple angles $\gamma_{\mathrm{P}}$, denoted as $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$, and for the radial polariser, denoted as $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$. The noise $\hat{\sigma}_{\mathrm{ext}}$ refers, in general, to the value specific to the data of each $\Lambda$, polarisation, and field-of-view (see Sec. III.1 for $\Lambda$ and $\gamma_{\mathrm{P}}$ values used). The wavelength dependence of the noise is indicated by the white triangles in the inset to Fig. 2.
@@
where $w(\lambda)=l(\lambda)c(\lambda)$ is the weighting incorporating the normalized illumination spectral intensity, $l(\lambda)$, and camera sensitivity, $c(\lambda)$, as functions of wavelength $\lambda$. The lamp spectrum is modeled as a blackbody radiaton at temperature, T=2800T=2800\,K, as described in Section 2.1.4 of Ref. 54. The camera sensitivity is known from the manufacturer and was digitized for this work. The wavelength dependence of $w(\lambda)$ is shown in Fig. LABEL:SM-fig:weights.
Figure 2: Measured properties of $N=51$ UGNSs of $D=30\,\mathrm{nm}$ nominal diameter. a) Asymmetry $\alpha_{\Lambda}$ versus cross-section $\sigma_{\Lambda}$ for $\Lambda=\{{\color[rgb]{0,0,1}450\,\mathrm{nm}},\,{\color[rgb]{0,0.92,0.92}500\,\mathrm{nm}},\,{\color[rgb]{0,0.9,0}550\,\mathrm{nm}},\,{\color[rgb]{1,0.6,0}600\,\mathrm{nm}},\,{\color[rgb]{1,0,0}650\,\mathrm{nm}},\,{\color[rgb]{0.7,0,0}700\,\mathrm{nm}}\}$, color-coded to reflect the center wavelength. Coloured crosses indicate the fitted parameters for linear polarisations using Eq. 9. Coloured vertical bars indicate the measured $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ for radial polarisation, at vertical positions chosen for clarity. A representative TEM image[43, 44] of NPs from the same batch is shown on the right. The noise in $\alpha_{\Lambda}$ is shown as shaded areas for $\Lambda={\color[rgb]{0,0,1}450\,\mathrm{nm}}$ (bluish) and $\Lambda={\color[rgb]{0.7,0,0}700\,\mathrm{nm}}$ (reddish), estimated as $(\hat{\sigma}_{\mathrm{ext}}/\sqrt{1.5})/\sigma_{\Lambda}$, corresponding to the lowest and highest noise versus $\Lambda$, respectively. See Sec. LABEL:SM-sec:optnoise for further details about the noise in the fit parameters. Solid vertical lines indicate the calculated $\sigma_{\Lambda}$ for a spherical gold NP for each $\Lambda$, taking into account averaging over filter ranges, for the sample-averaged TEM diameter, $27.94\,\mathrm{nm}$, and the OL permittivity dataset with $g=1.8$. The corresponding spectrum is given in the inset, showing additionally the case $g=0$. Dotted vertical lines indicate the same as solid, but calculated using Mie theory. The corresponding spectrum is shown by the dotted curve in the inset. Vertical colour bands indicate the filter ranges used in the experiment. The measurement error $\hat{\sigma}_{\mathrm{ext}}$, polarisation-averaged and FOV-averaged, are shown for each $\Lambda$, scaled by a factor $100$. b) Histograms of $\delta_{\sigma}$, for each $\Lambda$ as labelled, with gray bands showing $\pm\sqrt{2}\,\hat{\sigma}_{\mathrm{ext}}$, where the $\sqrt{2}$ factor accounts for the additional noise arising from the subtraction in $\delta_{\sigma}$.

### v2

@@ Abstract
Quantitative polarisation-resolved optical extinction microscopy of individual plasmonic nanoparticles has recently been introduced as a powerful tool to characterise the nanoparticle’s morphology with a precision comparable to electron microscopy, while using a simple optical microscope [Nanoscale 12, 16215 (2020)]. Here we provide a step change by adding measurements for radial polarisation in the condenser back focal plane, probing plasmonic resonances polarised in axial direction. The combined linear and radial polarisation measurements provide a significantly enhanced precision of the retrieved 3D morphology, as we show on defect-free ultra-uniform gold nanospheres of 30 nm nominal diameter characterised by transmission electron microscopy. The measured cross-sections are modelled for an ellipsoidal particle, determining the three semi-axes and rotation angles by fitting the measurements. The material permittivity and surface damping providing the best fit are found. The particle aspect ratio is determined with a precision better than 5%, and the size with an impressive precision of $0.1\,\mathrm{nm}$. Notably, corrections to the Rayleigh-Gans ellipsoid model due to retardation are significant even though the particle diameters are more than an order of magnitude smaller than the wavelength in the medium. Taking them into account improves the shape accuracy.
@@ I Introduction
Recently, our group established a methodology for 3D morphometry of single NPs called the Optical Nanosizer ($\mathrm{ON}$).[43] As a wide-field microscopy modality, the ON offers simultaneous measurement of up to hundreds of NPs in a single field of view, with a sensitivity down to $\sigma_{\mathrm{ext}}\approx 1\,$nm2, similar to the cross-section of a gold nanosphere (GNS) of $D\approx 2\,$nm diameter. In this method, a series of brightfield transmission images of NPs deposited on a coverslip are captured on a conventional microscope under illumination at different wavelengths and linear polarisations. Dark-field images can be additionally taken to distinguish scattering and absorption cross-sections.[35] Images are analysed as differential transmission contrast, to quantify $\sigma_{\mathrm{ext}}$ for each NP in the field of view. By comparing the Rayleigh–Gans model for ellipsoidal particles with the measurements of $\sigma_{\mathrm{ext}}$ for each NP in the images, we inferred the size and shape of nominally spherical small (10 to 30 nm diameter) gold nanoparticles, and gold nanorods (10 nm diameter, 30 nm length). The precision of the method was found to be $\pm 0.5\,\mathrm{nm}$ for size, and about $10\%$ for the shape anisotropy. In terms of accuracy, we found the measurements of GNS having nominal diameter, D=30D=30\,nm, to be within 11\,nm of TEM characterization on the same sample, i.e. showing only about $4\%$ discrepancy. Notably, for shape retrieval we used linearly polarised excitation light in the condenser back focal plane with varying in-plane angle. Even with a high numerical aperture (NA) condenser focussing the light, only a small fraction of light at the sample is polarised out of plane, i.e. axially. Hence, the measurements are dominated by excitation polarisations in the sample plane, and thus are more sensitive to the in-plane geometry of the NP. The out-of-plane geometry was therefore indirectly retrieved by spectral properties using the dependence of surface plasmon resonance frequencies on particle morphology.
@@ II Model: Permittivity, polarisability, and optical cross-sections
The previous $\mathrm{ON}$ analytical developments were described in detail in Ref. 43, 44. We build here on this analysis, extending it to add the treatment of the radially polarised illumination. We briefly recap the model in the following. The morphometry of each particle is derived from the measured cross-sections by solving ‘the inverse problem,’ using a theoretical framework relating the geometry of a NP to its optical cross-sections. Modelling the optical cross-sections requires knowledge of the complex permittivity of the NP material. In the literature, different measured spectra of the permittivity of gold are provided. We considered three of these, namely the data of Johnson & Christy (JC),[45] McPeak et al. (MP),[46] and Olmon et al. (OL).[47] They were measured by spectroscopic ellipsometry on thin polycrystalline films (JC, MP) or single crystalline gold (OL). To account for a surface damping effect in the NPs, we use the following approach. The permittivity is described by a Drude model with a bound electron contribution stemming from the interband transitions from the $d$ bands into the conduction band, and surface damping effects, given by,[48, 35]
@@
We use the Rayleigh–Gans model, describing ellipsoidal NPs much smaller than the light wavelength ($\lambda$), to determine the NP polarisability using the complex material permittivity and geometry of the NP, and the permittivity, $\epsilon_{\mathrm{m}}$, of the surrounding medium. As intruduced in Ref. 44, we use a Cartesian reference system of unit vectors $\vec{e}_{\kappa}^{\prime}$, with $\kappa=x^{\prime},\,y^{\prime}$, and $z^{\prime}$ which respectively point in the directions of the three semi-axes of an ellipsoidal NP of length ($a$, $b$, $c$), which we order such that $a\geq b\geq c$. In this basis, the polarisability tensor $\alpha^{\prime}$ is a diagonal matrix. The diagonal elements are the polarisabilities for fields oriented along each of the NP’s semi-axes, given by
@@
where $V$ is the NP volume, $\epsilon_{0}$ is the free space permittivity, $\epsilon_{\mathrm{m}}$ is the dielectric function of the surrounding medium, and $L_{j}$ are depolarisation factors, with $j\in\{1,\,2,\,3\}$. The geometry of the NP determines $L_{j}$. For example, for a sphere, $L_{j}=1/3$. $\epsilon_{\mathrm{m}}$ is assumed to be real and frequency independent. All materials are assumed to be non-magnetic, i.e. with relative permeability $\mu_{\mathrm{r}}=1$. We note that metals exhibit a zero crossing versus wavelength of the real part of the denominator in Eq. 3 at $\Re\{\epsilon_{\mathrm{NP}}\}=(1-L_{j}^{-1})\epsilon_{\mathrm{m}}$, which results in a peak in $\sigma_{\mathrm{ext}}$ called the localised surface plasmon resonance (LSPR). For an incident field $\vec{E}^{\prime}$, the induced electric dipole moment is given by
@@
We choose a Cartesian coordinate system to describe the laboratory frame, with axes $\iota\in\{x,y,z\}$, and associated unit vectors $\vec{e}_{\iota}$, where $\vec{e}_{z}$ points along the optical path, and $\vec{e}_{x}$, $\vec{e}_{y}$ span the sample plane. To transform $\alpha^{\prime}$ into the laboratory reference frame, we define[43] a 3D rotation matrix using $R=R_{\psi}R_{\theta}R_{\phi}$ and $R^{\intercal}=R_{\phi}^{\intercal}R_{\theta}^{\intercal}R_{\psi}^{\intercal}$, with $\phi$, $\theta$, $\psi$ the angles of rotation about $\vec{e}_{x}$, $\vec{e}_{y}$, $\vec{e}_{z}$, respectively. We transform a vector from the NP frame into the laboratory frame with $\vec{v}=R\vec{v}^{\prime}$. Eq. 4 becomes $R^{\intercal}\vec{p}=\alpha^{\prime}R^{\intercal}\vec{E}$, such that the polarisability in the laboratory frame is given by
@@
To compare the ellipsoid geometry with plan-view TEM characterisation, we define a mean aspect ratio (MAR), $\sqrt{bc}/a$, and a projected aspect ratio (PAR) given by $b^{\prime}/a^{\prime}$, with the semi-axes lengths $a^{\prime}$ and $b^{\prime}$ in the $xy$-plane of the ellipsoid projected along the $z$ axis (for calculation details see Sec. B). We also define the corresponding projected diameter as $D_{\mathrm{p}}=2\sqrt{a^{\prime}b^{\prime}}$ and a volume-equivalent diameter as $D_{V}=2\sqrt[3]{abc}$.
@@ III.3 Optical Extinction Measurements
In Ref. 53 we detailed our procedure to quantitatively measure $\sigma_{\mathrm{ext}}$, with additional information given in Ref. 54. In short, two brightfield images called $I_{1}$ and $I_{2}$ are captured, which differ only by a lateral shift of the sample, typically 1 to 2 µm (i.e. a few optical resolutions). $I_{1}$ and $I_{2}$ are averaged in real time over a number, $N_{\mathrm{i}}$, of individual acquisitions to reduce shot noise. We compute $\Delta_{1,2}=1-I_{1,2}/I_{2,1}$ to yield two extinction images (with inverted contrast). To further reduce shot noise as well as systematic noise due to sensor electronic drift, we can average the extinction over a number of repetitions $N_{\mathrm{r}}$. For a NP centered in the areas $A_{1,2}$ of radius $r_{\mathrm{i}}$ in $I_{1,2}$, we determine its extinction as $2\sigma_{\mathrm{ext}}=\int_{A_{1}}\Delta_{1}\mathrm{d}A+\int_{A_{2}}\Delta_{2}\mathrm{d}A$. We calculate the shot-noise-limited standard deviation in the measured $\sigma_{\mathrm{ext}}$ as [53]
@@
where $M$ is the magnification from the sample to the detector, $N_{\mathrm{a}}=N_{\mathrm{i}}N_{\mathrm{r}}$ is the number of acquired frames, and $d_{\mathrm{px}}=6.5\,$µm is the pixel pitch of the sensor. We previously used[34] a measurement radius of $r_{\mathrm{i}}=3\lambda/(2\mathrm{NA})$ to analyse $\sigma_{\mathrm{ext}}$, and the effect of $r_{\mathrm{i}}$ on $\sigma_{\mathrm{ext}}$ is characterised in detail in Ref. 55. In the present work, we employ three values of $r_{\mathrm{i}}$, specifically $r_{\mathrm{s}}=3\lambda/(2\mathrm{NA})$, $r_{\mathrm{h}}=1.2\,r_{\mathrm{s}}$, and $r_{\mathrm{l}}=0.8\,r_{\mathrm{s}}$, with $N_{\mathrm{i}}=128$ and $N_{\mathrm{r}}=20$. At $\Lambda=550\,$nm, Eq. 8 yields $\hat{\sigma}_{\mathrm{ext}}\approx 7.5\,\mathrm{nm}^{2}$ using $r_{\mathrm{s}}$, while the noise seen in the extinction data at this wavelength is $15.6\,\mathrm{nm}^{2}$. The additional noise is attributed to surface roughness, debris, and/or residual sensor fluctuations, as analysed in detail in Ref. 53. We note that measurements are taken over several different fields of view (FOV) and aggregated for the final statistics. The $\hat{\sigma}_{\mathrm{ext}}$ reported are the polarisation-averaged values of the noise, averaged over the measured values from all the FOVs. We found that $\hat{\sigma}_{\mathrm{ext}}$ varies by $\pm 5\%$ across the illumination polarisations, and by $\pm 15\%$ across FOVs. We measure $\sigma_{\mathrm{ext}}$ for multiple wavelengths $\Lambda$, both for the linear polariser at multiple angles $\gamma_{\mathrm{P}}$, denoted as $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$, and for the radial polariser, denoted as $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$. The noise $\hat{\sigma}_{\mathrm{ext}}$ refers, in general, to the value specific to the data of each $\Lambda$, polarisation, and field-of-view (see Sec. III.1 for $\Lambda$ and $\gamma_{\mathrm{P}}$ values used). The wavelength dependence of the noise is shown in the inset of Fig. 2a.
@@
where $w(\lambda)=l(\lambda)c(\lambda)$ is the weighting incorporating the normalized illumination spectral intensity, $l(\lambda)$, and camera sensitivity, $c(\lambda)$, as functions of wavelength $\lambda$. The lamp spectrum is modeled as a blackbody radiator at temperature, T=2800T=2800\,K, as described in Section 2.1.4 of Ref. 54. The camera sensitivity is known from the manufacturer and was digitized for this work. The wavelength dependence of $w(\lambda)$ is shown in Fig. 6.
Figure 2: Measured properties of $N=51$ UGNSs of $D=30\,\mathrm{nm}$ nominal diameter. a) Asymmetry $\alpha_{\Lambda}$ versus cross-section $\sigma_{\Lambda}$ for $\Lambda=\{{\color[rgb]{0,0,1}450\,\mathrm{nm}},\,{\color[rgb]{0,0.92,0.92}500\,\mathrm{nm}},\,{\color[rgb]{0,0.9,0}550\,\mathrm{nm}},\,{\color[rgb]{1,0.6,0}600\,\mathrm{nm}},\,{\color[rgb]{1,0,0}650\,\mathrm{nm}},\,{\color[rgb]{0.7,0,0}700\,\mathrm{nm}}\}$, color-coded to reflect the center wavelength. Coloured crosses indicate the fitted parameters for linear polarisations using Eq. 9. Coloured vertical bars indicate the measured $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ for radial polarisation, at vertical positions chosen for clarity. A representative TEM image[43, 44] of NPs from the same batch is shown on the right. The noise in $\alpha_{\Lambda}$ is shown as shaded areas for $\Lambda={\color[rgb]{0,0,1}450\,\mathrm{nm}}$ (bluish) and $\Lambda={\color[rgb]{0.7,0,0}700\,\mathrm{nm}}$ (reddish), estimated as $(\hat{\sigma}_{\mathrm{ext}}/\sqrt{1.5})/\sigma_{\Lambda}$, corresponding to the lowest and highest noise versus $\Lambda$, respectively. See Sec. C for further details about the noise in the fit parameters. Solid vertical lines indicate the calculated $\sigma_{\Lambda}$ for a spherical gold NP for each $\Lambda$, taking into account averaging over filter ranges, for the sample-averaged TEM diameter $D_{V}$, $27.94\,\mathrm{nm}$, and the OL permittivity dataset with $g=1.8$. The corresponding spectrum is given in the inset, showing additionally the case $g=0$. Dotted vertical lines indicate the same as solid, but calculated using Mie theory. Dash-dotted lines indicate the same as dotted, but for MP permittivity. The corresponding spectrum is shown by the dotted curve in the inset. Vertical colour bands indicate the filter ranges used in the experiment. The measurement error $\hat{\sigma}_{\mathrm{ext}}$, polarisation-averaged and FOV-averaged, are shown for each $\Lambda$, scaled by a factor $100$. b) Histograms of $\delta_{\sigma}$, for each $\Lambda$ as labelled, with gray bands showing $\pm\sqrt{2}\,\hat{\sigma}_{\mathrm{ext}}$, where the $\sqrt{2}$ factor accounts for the additional noise arising from the subtraction in $\delta_{\sigma}$.